OpenStudy (anonymous):

A vector equation being a part of line L is expressed as something I'll show below. How do I know which vector equations are on the same line L? *see details*

OpenStudy (anonymous):

r = \[\left(\begin{matrix}4 \\ 4\end{matrix}\right)\] + \[\left(\begin{matrix}3 \\ 1\end{matrix}\right)\]

OpenStudy (anonymous):

and what does (4 (4 below 4 here) ) mean for example? does it have something to do with i j and k?

OpenStudy (theeric):

If two vectors are in line with one another, then I think the magnitude of the cross product of the two vectors should be 0.